OpenStudy (anonymous):

integrate sqrt(25-x^2)/x^2

OpenStudy (anonymous):

i've worked through it and would like some feedback on my steps

OpenStudy (anonymous):

\[\frac{ \sqrt{25-x^{2}} }{ x ^{2} }\]

OpenStudy (anonymous):

let x=5sin theta dx=5cos theta

OpenStudy (anonymous):

\[\frac{ \sqrt{5^2-5^2\sin \theta} }{ 5^2\sin \theta } 5\cos \theta\]